Two rods of diffrent material with diffrent area of cross section and length of second rod is equal to twice of length of first rod, k2=4kand diameter of d2=d1/2 where d is the diameter  if two conductors . Find the Temperature at the junction 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

For rods connected end to end, the rate of heat transfer through them must be same.

Hence\dfrac{K_1A_1}{l_1}(T_A-T)=\dfrac{K_2A_2}{l_2}(T-T_B)
\implies \dfrac{K_1\pi d_1^2}{4l_1}(T_A-T)=\dfrac{K_2\pi d_2^2}{4l_2}(T-T_B)
\implies 2(T_A-T)=(T-T_B)
\implies T=\dfrac{2T_A+T_B}{3}